a.   Calculate the magnetic induction that exists in air at a distance of 20 mm from a long wire that conducts a current of 6.5A.

b.   Find the magnetic induction B that exists in air at a point located at 15 cm from a long wire that conducts a current of 2A.



1
Expert's answer
2021-11-21T17:27:16-0500

a.


"B=\mu_0I/(2\pi r)=4\pi\cdot10^{-7}\cdot 6.5/(2\pi\cdot0.02)="


"=6.5\cdot 10^{-5}\ (T)"


b.


"B=\mu_0I/(2\pi r)=4\pi\cdot10^{-7}\cdot 2/(2\pi\cdot0.15)="


"=2.7\cdot 10^{-6}\ (T)"